A parallel beam of red light of wavelength 650nm arrives normally at a diffraction grating. The first order maximum has an angle of 40°. Calculate the total number of principal maxima this grating can produce for the red light. At what angle is the highest order maxima produced?


1
Expert's answer
2021-03-03T07:56:25-0500

Answer

Relationship is given

the total number of principal maxima this grating can produce for the red light.

"N=\frac{sin\theta}{n\lambda}\\=\frac{sin40^°}{1\times650\times10^{-9}}\\=989\times10^{3}"


Now


At what angle is the highest order maxima produced is "\theta=90^°"